OpenStudy (anonymous):

Variable g is 8 more than variable w. Variable g is also 4 less than w. Which pair of equations best models the relationship between g and w? A. g= w – 8 g = w + 4 B. w = 8g w = g – 4 C. g= w + 8 g = w – 4 D. w = 8g w = g + 4

Vocaloid (vocaloid):

"8 more than w" means that g = w + 8 (more means add, in this case) "4 less" means that g = w - 4 (less means subtract, in this case)

OpenStudy (hockeychick23):

g=w+8 g=w-4